Verify that a linear transformation T from R2 into R2    

Q1: Verify that a linear transformation T from RP into R T(V1, V2) = (v1 – V2, V1 + 2v2)
Answer #1

Given: TR² R² is defined as TLV, Vz) = (V-V2, Y +242) & Let u =(4,4), V = CV, 2) ER² Then, T(F) = T(4,4) =(4,-4, 4, +242) T(E